A car is traveling on a road that is perpendicular to a railroad track. When the car is 35 meters from the crossing, the car's new collision detector warns the friver that there is a train 60 meters from the car and heading toward the same crossing.How far (in meters) is the train from the crossing?
Answer by rfer(16322)

x^2+35^2=60^2
x^2+1225=3600
x^2=3600-1225
x^2=2375
x=48.7 m